consider a papulation consisting 2,5,8,11,13 and 25. Suppose sample of size 5 are drawn from this population.

1
Expert's answer
2022-03-07T04:15:03-0500

We have population values "2,5,8,11,13,25" population size "N=6" and sample size "n=5."

Thus, the number of possible samples which can be drawn without replacements is "\dbinom{6}{5}=6."

1.


"mean=\mu=\dfrac{2+5+8+11+13+25}{6}=\dfrac{32}{3}"

2


"Variance=\sigma^2=\dfrac{1}{6}((2-\dfrac{32}{3})^2+(5-\dfrac{32}{3})^2"




"+(8-\dfrac{32}{3})^2+(11-\dfrac{32}{3})^2+(13-\dfrac{32}{3})^2"




"+(25-\dfrac{32}{3})^2)=\dfrac{488}{9}"


"\sigma=\sqrt{\sigma^2}=\sqrt{\dfrac{488}{9}}=\dfrac{2\sqrt{122}}{3}\approx7.363574"



3.


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 Sample\ values & Sample\ mean\ (\bar{x}) \\ \hline
 2,5,8,11,13 & 7.8 \\
 \hdashline
 2,5,8,11,25 & 10.2 \\
 \hdashline
 2,5,8,13,25 & 10.6 \\
 \hdashline
 2,5,11,13,25 & 11.2\\
 \hdashline
 2,8,11,13,25 & 11.8 \\
 \hdashline
 5,8,11,13,25 & 12.4 \\
 \hdashline
\end{array}"



5. The sampling distribution of the sample mean "\bar{x}" and its mean and standard deviation are:


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 & \bar{x} & f & f(\bar{x}) & \bar{x}f(\bar{x}) & \bar{x}^2f(\bar{x})\\ \hline
 & 7.8 & 1 & 1/6 & 7.8/6 & 60.84/6 \\
 \hdashline
 & 10.2 & 1 & 1/6 & 10.2/6 & 104.04/6 \\
 \hdashline
 & 10.6 & 1 & 1/6 & 10.6/6 & 112.36/6 \\
 \hdashline
 & 11.2 & 1 & 1/6 & 11.2/6 & 125.44/6 \\
 \hdashline
 & 11.8 & 1 & 1/6 & 11.8/6 & 139.24/6 \\
 \hdashline
 & 12.4 & 1 & 1/6 & 12.4/6 & 153.76/6 \\
 \hdashline
 Sum= & & 6 & 1 & 32/3 & 695.68/6 \\
 \hdashline
\end{array}"


6.


"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{x}f(\bar{x})=\dfrac{32}{3}"



7.


"Var(\bar{X})=\sigma_{\bar{X}}^2=\sum\bar{x}^2f(\bar{x})-(\sum\bar{x}f(\bar{x}))^2"




"=\dfrac{347.84}{3}-(\dfrac{32}{3})^2=\dfrac{19.52}{9}"




"\sigma_{\bar{X}}=\sqrt{\sigma_{\bar{X}}^2}=\sqrt{\dfrac{19.52}{9}}\approx1.472715"



Check


"\mu_{\bar{X}}=E(\bar{X})=\dfrac{32}{3}=\mu"




"Var(\bar{X})=\dfrac{19.52}{9}=\dfrac{\sigma^2}{n}\cdot\dfrac{N-n}{N-1}=\dfrac{488}{9(5)}\cdot\dfrac{6-5}{6-1}"